[figure: 9.7 cm]

MONTE CARLO ANALYSIS OF THE ASHKIN-TELLER MODELS

L. Dȩbski Computational Physics Division, Institute od Physics A. Mickiewicz University, ul. Umultowska 85, 61-614 Poznan, Poland

Monte Carlo simulations are used to analyse the critical behaviour of the Ashkin-Teller model [1] in two and three dimensions from the ratio of the square of the second moment of the order parameter to its fourth moment [2]. Phase diagram is calculated in regions of the Ising transitions. The continuity of the order parameter at the critical surface is verified by a finite size scaling analysis for both ferromagnetic and antiferromagnetic interactions.

[1] R.J. Baxter, Exactly Solvable Models in Statistical Mechanics, Academic Press, London, 1982

[2] P. Pawlicki, G. Kamieniarz and L. Dȩbski, Physica A242, 290 (1997)

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : L. Dȩbski

Address for correspondence : Computational Physics Division, Institute od Physics A. Mickiewicz University, ul. Umultowska 85, 61-614 Poznan, Poland

Email address : lech@pearl.amu.edu.pl